Help needed please.
Steve traveled 600 miles at a certain speed. Had he gone 20mph faster, the trip would have taken 1 hour less. Find the speed of his vehicle.

Answer by checkley75(3666)

600/x=600/(x+20)+1
600/x=(600+x+20)/(x+20)
x(600+x+20)=600(x+20)
x(620+x)=600x+1200
620x+x^2-600x-1200=0
x^2+20x-1200=0
(x-100)(x+120)=0
x-100=0
x=100 answer
proof
600/100=600/120+1
6=5+1
6=6